OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (anonymous):

since line is parallel to y=3x-6 so the slope of required line is 3. So equation of line y+4=3(x-2) y=3x-10